Why is $10 to $20 a 100% increase, while $20 to $10 is a 50% decrease. Explain your answer.
Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

Why is $10 to $20 a 100% increase, while $20 to $10 is a 50% decrease. Explain your answer.
--------------
It's based on what the change is from.
From 10 to 20 starts with 10. 10/10 is 100%
From 20 to 10 starts with 20. 10/20 is 50%
-----------------
What you start with is the base.
If you try to do it any other way, it's a lot less meaningful.

Answer by richwmiller(17219)

The increase from 10 to 20 is 10
10/10=1 = 100%
but decrease from 20 to 10 is 20-10=10
10/20=.5=50%